### v3

@@ Abstract
This work addresses the pressing need for cybersecurity in Unmanned Aerial Vehicles (UAVs), particularly focusing on the challenges of identifying UAVs using radiofrequency (RF) fingerprinting in constrained environments. The complexity and variability of RF signals, influenced by environmental interference and hardware imperfections, often render traditional RF-based identification methods ineffective. To address these complications, the study introduces the rigorous use of one-shot generative methods for augmenting transformed RF signals, offering a significant improvement in UAV identification. This approach shows promise in low-data regimes, outperforming deep generative methods like conditional generative adversarial networks (GANs) and variational auto-encoders (VAEs). The paper provides a theoretical guarantee for the effectiveness of one-shot generative models in augmenting limited data, setting a precedent for their application in limited RF environments. This research not only contributes to the cybersecurity of UAVs but also rigorously broadens the scope of machine learning techniques in data-constrained scenarios, which may include atypical complex sequences beyond images and videos.
{IEEEkeywords}
Low-Data Regimes, Wasserstein Metric, Synthetic Data, Out-of-distirbution, Extrapolation.
@@ 1 INTRODUCTION
[…]
The growing integration of Unmanned Aerial Vehicles (UAVs), commonly known as drones, into various sectors, ranging from surveillance to logistics, has underscored the critical need for their cybersecurity. Given the risk of cyber-attacks on UAVs, employing radiofrequency (RF) fingerprinting becomes crucial for identifying legitimate versus malicious devices, particularly in environments with constraints [1, 2]. The complexity and variability of RF signals, influenced by factors such as temperature, humidity [3], environmental obstructions [4, 5], and hardware imperfections [6], pose significant challenges for consistent UAV identification and tracking. Traditional RF-based identification methods, although effective in open environments, are often compromised in areas with dense foliage, urban canyons, or inclement weather conditions, leading to unreliable UAV detection and classification.
The advent of machine learning techniques has opened new avenues for enhancing UAV identification in these challenging RF environments. By leveraging deep learning algorithms, it is possible to extract and analyze complex patterns within the RF signals that are typically obscured or distorted due to environmental factors [7, 8, 9, 10]. However, the application of data-intensive deep learning methods does not inherently resolve the issue of data scarcity in compromised environments. This necessitates the development of reliable techniques for augmenting RF signals or their transformations in situations with limited data. The challenge arises mainly because current data augmentation techniques are predominantly tailored for computer vision applications, and their extension to other types of sequences may not yield the anticipated outcomes due to the complex semantics of the signals.
In this study, we rigorously employ a one-shot generative method, inspired by computer vision techniques, for the augmentation of transformed RF signals across various training data sizes. The results demonstrate a significant improvement in evaluation metrics for the UAV identification, outperforming other frameworks such as conditional generative adversarial networks (GAN) and variational auto-encoders (VAE) in addressing low-data regime circumstances  [^1] https://github.com/amir-kazemi/UAV_ID_limited_RF.
The work is therefore structured as follows: First, we investigate the ability of one-shot generative models for augmenting a dataset of sequences. Particularly, we provide the mathematical definition of sequence, subsequence, one-shot generative models, and distance metrics. We establish an upper bound for the Wasserstein distance between the real and generated sequences, considering the expected distances of their subsequences. We also validate our general guarantee by intuitive special cases of sequences and subsequences. Secondly, we go through the UAV identification experiment and elaborate on the selected dataset, the one-shot generative model, benchmark conditional deep generators, the classification pipeline, and the evaluation method. Finally, we discuss the results which include the accuracy, precision, recall, and F1-score of the classification task for all designed cases and methods.
@@ 2 RELATED GENERATIVE TECHNIQUES
Synthetic data generation is a crucial field of study in machine learning and data science, with potential applications spanning a diverse array of disciplines. In many cases, obtaining large amounts of high-quality, labeled data is a significant challenge, and this scarcity can lead to biased or skewed results when training models [11]. Data generation methods can help create edge cases, rare events, and other scenarios that are difficult to collect in real-world circumstances, allowing for more robust model training [12].
The past decade marked a dramatic increase in efforts to generate sequences for a broad range of data, from images and videos to audio, text, and spatiotemporal physical states. This surge in sequence generation can be attributed largely to the recent advances in convolutional and recurrent neural networks (CNN and RNN) [13, 14, 15, 16] and deep generative frameworks such as GAN [17] and VAE [18]. At the heart of these architectures and frameworks lie deep neural networks (DNN). However, when training generalizable DNNs using Empirical Risk Minimization (ERM) [19], it is paradoxical to depend on deep generative models that require large amounts of data to address data scarcity in subsequent tasks. This situation highlights the necessity to explore alternative approaches that can address the challenges of data-intensive models, utilizing a concept known as Vicinal Risk Minimization (VRM).
VRM is a training principle that improves generalization in machine learning models by considering the vicinity of the training samples, hypothesizing that slight perturbations around real data points can yield more robust and better-performing models [20]. This concept seamlessly leads to the notion of data augmentation, a widely-adopted strategy in deep learning, especially in fields like computer vision where common methods of data augmentation include random cropping, rotation, and brightness adjustments [21].
With the emergence of one-shot generative models, a fresh motivation surfaces; harnessing their potential to generate realistic variations of training samples might offer a more sophisticated approach to VRM. Leveraging such generative capacities could potentially revolutionize the landscape of data augmentation and VRM practices. Recent one-shot generative models [22, 23, 24, 25], and their non-deep successors [26, 27, 28], generate patch-similar images and videos from a single target. Considering the patch-similarity as a vicinity criterion, one may therefore explore the ability of one-shot generative models for augmenting an entire dataset inspired by the VRM principle.
In this study, our focus is on the application of one-shot generative models to sequences, which serve as a universal structure for ordered data, extending from images to time series. We aim to establish theoretical limits on their ability to augment datasets. Our work utilizes the Generating by Patch Distribution Matching (GPDM) algorithm, detailed in [27]. GPDM is designed to minimize the Wasserstein distance between patches of both real and generated images. Furthermore, this algorithm employs a multi-scale hierarchical strategy, facilitating an early understanding of the global image structure during the initial training stages.
@@ 3 SEQUENCE AND SUBSEQUENCE SIMILARITY
According to the definition, $\Omega_{S}$ is the set of all matrices created by choosing $d^{\prime}$ rows from the identity matrix $\mathbf{I}_{d}$. The probability distribution for singleton events in $\tilde{\Omega}_{S}\subseteq\Omega_{S}$ is uniform. Furthermore, the subset $\tilde{\Omega}_{S}$ must be sufficiently large to allow for the recovery of the sequence $\mathbf{X}$ from the projections $(\mathbf{s}^{T}\mathbf{s}\mathbf{X})$, as per (2). This suggests that $\bm{\lambda}\in\mathbb{N}^{d}$, since $\bm{\lambda}$ represents the frequency of occurrences of elements from $\mathbf{X}$ in these projections, as illustrated in Fig. 1.
Based on this definition, we can proceed to define the behavior of a one-shot generative model. This model is designed to produce subsequence-similar sequences for every target sequence within the dataset. Before proceeding, we elucidate our understanding of similarity by referencing a conventional metric distance used to compare probability distributions.
@@
$\begin{split}&\|\mathbf{X}^{\prime}-G(\mathbf{X},\mathbf{Z})\|\\ \leq&\textstyle\bm{\lambda}_{\min}^{-1}\norm{\sum_{\mathbf{s}\in\tilde{\Omega}_{S}}(\mathbf{s}^{T}\mathbf{sX^{\prime}}-\mathbf{s}^{T}\mathbf{s}G(\mathbf{X},\mathbf{Z}))}\\ \leq&\textstyle\bm{\lambda}_{\min}^{-1}\sum_{\mathbf{s}\in\tilde{\Omega}_{S}}\norm{\mathbf{s}^{T}\mathbf{sX^{\prime}}-\mathbf{s}^{T}\mathbf{s}G(\mathbf{X},\mathbf{Z})}\\ =&\textstyle|\tilde{\Omega}_{S}|\bm{\lambda}_{\min}^{-1}\mathbb{E}_{S}\norm{\mathbf{S}^{T}\mathbf{SX^{\prime}}-\mathbf{S}^{T}\mathbf{S}G(\mathbf{X},\mathbf{Z})}\end{split}$ (14)
@@
Assume that $\forall\delta>0$, there exists a measurable function $G$ such that: $\mathbb{E}_{X}\mathbb{E}_{Z}\inf\mathbb{E}_{S,S^{\prime}}\norm{\mathbf{S}\mathbf{X}-\mathbf{S}^{\prime}G(\mathbf{X},\mathbf{Z})}\leq\delta$. Let $\mathbf{X}=\textup{vec}(\mathbf{Y})$ where $\mathbf{Y}$ belongs to $(\Omega_{Y}\subset\mathbb{R}^{n\times n},P_{Y})$. Then, if $\tilde{\Omega}_{S}$ denotes the set of all $n^{\prime}\times n^{\prime}$ substrings where $n^{\prime}\leq n$, we have follwoing bounds:
@@
4 EXPERIMENTS
@@ 4.1 DATASET
RF sensing techniques for UAV detection and identification are practical for real-world use. These methods do not depend on UAVs’ wireless protocols like Bluetooth, 4G, or WiFi, and can manage physical differences across UAV models. This study utilizes a UAV RF dataset [7, 29] which encompasses RF signals obtained from three distinct UAV variants across a range of flight scenarios and is readily available for public access [^2] https://github.com/Al-Sad/DroneRF. To facilitate the objectives of UAVs detection and identification, the authors employed Fourier transform to process the RF signals, thereby extracting their frequency components. Specifically, the dataset supports three primary tasks: the detection of a UAV’s presence (two classes), the detection of a UAV’s presence coupled with the identification of its type (four classes), and the detection of a UAV’s presence alongside the identification of its type and flight mode (ten classes). In this work, we skip the detection task as it may not be a challenge even for limited RF environments.
RF signals are stored as segments to prevent memory overflows, and each segment is then processed using Discrete Fourier Transform (DFT). This method is advantageous for classification tasks for several reasons. First, it reveals important characteristics often hidden in the time domain by focusing on frequency components. Second, DFT reduces the dimensionality of the data, which is particularly beneficial in applications with high-dimensional or noisy raw data. Third, in applications like RF detection, where raw RF data may vary in length due to different UAV flight times, DFT ensures uniform length by concentrating on a specific frequency bandwidth for classification. Fourth, DFT facilitates the extraction of specific frequency features crucial to the classification or identification tasks. Fifth, frequency components are generally less impacted by noise than time-domain signals, rendering DFT-based features more robust against common issues like Electromagnetic Interference (EMI) and Radio Frequency Interference (RFI) in UAV RF signals. In addition, DFT-based features are often transferable across different signal processing tasks. Lastly, classifiers tend to perform better when trained on DFT-based features, as the inherent structure and discrimination power of frequency features can lead to improved accuracy and generalization.
The DFT of the unprocessed RF signals is computed as follows for lower and high frequency bands, respectively:
$y_{i}^{L}(m)=\textstyle\left|\sum_{n=1}^{N}x_{i}^{L}(n)\exp{-2\pi m(n-1)j\mathbin{/}N}\right|$ (27)
$y_{i}^{H}(m)=\textstyle\left|\sum_{n=1}^{N}x_{i}^{H}(n)\exp{-2\pi m(n-1)j\mathbin{/}N}\right|$ (28)
where $N$ signifies the raw RF signal’s length, $(x_{i},y_{i})$ represents the raw RF and spectrum of the $i$-th segment. The magnitude operator, denoted by ||$\cdot$|| ⋅ |, captures the power spectrum. Superscripts $L$ and $H$ differentiate between lower and higher frequency bands. This division is due to RF capture receivers’ technical limitations, requiring different devices for varying frequency ranges.
The complete RF spectrum of the $i$-th segment is formed by concatenating and normalizing spectra from both lower and higher frequency bands:
$y_{i}=\left[y_{i}^{L},cy_{i}^{H}\right]$ (29)
Here, $c=\sum_{q=0}^{Q}y_{i}^{L}(M-q)/\sum_{q=0}^{Q}y_{i}^{H}q$ is the normalization factor, ensuring spectral continuity between lower and higher band spectra. $M$ represents the total frequency indices in $y_{i}$. The selection of $Q$ is crucial for normalization. Lower $Q$ values facilitate smooth combination of the two spectra, while higher values help reduce random fluctuations, presenting a trade-off between seamless integration and fluctuation mitigation. In the context of the dataset under consideration, the specific parameter values chosen are $M=2048$ and $Q=10$.
In our application, we select the first 2025 features out of 2048 from the DFT-transformed RF signals and arrange them into a $45\times 45$ matrix structure for one-shot generation, as per Corollary 2. We also randomly select 1,135 out of 22,700 samples from the dataset to simulate low-data regime circumstances, ensuring data integrity while maintaining label proportions. Each sample in this dataset is linked to a label, which is a one-hot encoded vector of size four or ten, depending on the classification task, with the first class designated for background noise in no-UAV scenarios.
Initially, we apply a K-fold strategy (with $K=5$) to the original dataset, setting aside a consistent 20% random sample for testing and verification in each fold, with proportional representation of each label. This method ensures that these test samples remain unused in the training phase, preserving their utility for evaluating model performance against adversarial samples. The training phase then exclusively employs the remaining 80% of the dataset, totaling 908 samples in each fold which is called “complete” data throughout this work. For a comprehensive analysis, we further segment the training portion of each fold into reduced datasets, representing five scenarios with sizes between $10\sim 50\%$ of the training data. This segmentation allows us to demonstrate our model’s effectiveness across a range of reduced training dataset sizes. To ensure fair model performance comparisons, 5000 synthetic datasets are created from subsampled sets for all generative models. This method facilitates meaningful evaluations of models on synthetic versus original data across varying scenarios.

### v4

@@ Abstract
This work addresses the challenge of identifying Unmanned Aerial Vehicles (UAV) using radiofrequency (RF) fingerprinting in limited RF environments. The complexity and variability of RF signals, influenced by environmental interference and hardware imperfections, often render traditional RF-based identification methods ineffective. To address these complications, the study introduces the rigorous use of one-shot generative methods for augmenting transformed RF signals, offering a significant improvement in UAV identification. This approach shows promise in low-data regimes, outperforming deep generative methods like conditional generative adversarial networks (GANs) and variational auto-encoders (VAEs). The paper provides a theoretical guarantee for the effectiveness of one-shot generative models in augmenting limited data, setting a precedent for their application in limited RF environments. This research contributes to learning techniques in low-data regime scenarios, which may include atypical complex sequences beyond images and videos. The code and links to datasets used in this study are available at https://github.com/amir-kazemi/uav-rf-id.
keywords:
Low-Data Regimes, Wasserstein Metric , Synthetic Data , Out-of-distribution , Extrapolation.
One-shot Generative Distribution Matching for Augmented RF-based UAV Identification
[…]
University of Illinois at Urbana-Champaign, city=Urbana, postcode=61801, state=IL, country=USA
@@ 1 Introduction
The growing integration of Unmanned Aerial Vehicles (UAVs), commonly known as drones, into various sectors, ranging from surveillance to logistics, has underscored the critical need for their cybersecurity. As such, employing radiofrequency (RF) fingerprinting becomes crucial for identifying legitimate versus malicious devices, particularly in environments with constraints [1, 2]. The complexity and variability of RF signals, influenced by factors such as temperature, humidity [3], environmental obstructions [4, 5], and hardware imperfections [6], pose significant challenges for consistent UAV identification and tracking. These conditions can interfere with RF signal clarity and strength, leading to a limited RF environment where the number of usable signals is reduced. Traditional RF-based identification methods, although effective in open settings, often struggle in such limited RF environments—marked by dense foliage, urban canyons, adverse weather, or potentially exacerbated by hardware malfunctions—making reliable UAV detection and classification particularly challenging. The advent of machine learning techniques has opened new avenues for enhancing UAV identification in these challenging RF environments. By leveraging deep learning algorithms, it is possible to extract and analyze complex patterns within the RF signals that are typically obscured or distorted due to environmental factors [7, 8, 9, 10]. However, the application of data-intensive deep learning methods does not inherently resolve the issue of data scarcity in compromised environments. This necessitates the development of reliable techniques for augmenting RF signals or their transformations in situations with limited data. The challenge arises mainly because current data augmentation techniques are predominantly tailored for computer vision applications, and their extension to other types of sequences may not yield the anticipated outcomes due to the complex semantics of the signals. In this study, we rigorously employ a one-shot generative method, inspired by computer vision techniques, for the augmentation of DFT-transformed RF signals across various training data sizes which simulates low-data regimes posed by limited RF environments. The results demonstrate a significant improvement in evaluation metrics for the UAV identification, outperforming other frameworks such as conditional generative adversarial networks (GAN) and variational auto-encoders (VAE) in addressing low-data regime circumstances. The work is therefore structured as follows: First, we investigate the ability of one-shot generative models for augmenting a dataset of sequences. Particularly, we provide the mathematical definition of sequence, subsequence, one-shot generative models, and distance metrics. We establish an upper bound for the Wasserstein distance between the real and generated sequences, considering the expected distances of their subsequences. We also validate our general guarantee by intuitive special cases of sequences and subsequences. Secondly, we go through the UAV identification experiment and elaborate on the selected dataset, the one-shot generative model, benchmark conditional deep generators, the classification pipeline, and the evaluation method. Finally, we discuss the results which include the accuracy, precision, recall, and F1-score of the classification task for all designed cases and methods.
@@ 2 Related generative techniques
Synthetic data generation is a crucial field of study in machine learning and data science, with potential applications spanning a diverse array of disciplines. In many cases, obtaining large amounts of high-quality, labeled data is a significant challenge, and this scarcity can lead to biased or skewed results when training models [11]. Data generation methods can help create edge cases, rare events, and other scenarios that are difficult to collect in real-world circumstances, allowing for more robust model training [12]. Distribution matching, an integral part of synthetic data generation, involves adjusting the synthetic data’s statistical properties to closely resemble those of the real data, thereby enhancing the validity and effectiveness of the models trained on this data [13]. The past decade marked a dramatic increase in efforts to generate sequences for a broad range of data, from images and videos to audio, text, and spatiotemporal physical states. This surge in sequence generation can be attributed largely to the recent advances in convolutional and recurrent neural networks (CNN and RNN) [14, 15, 16, 17] and deep generative frameworks such as GAN [18] and VAE [19]. At the heart of these architectures and frameworks lie deep neural networks (DNN). However, when training generalizable DNNs using Empirical Risk Minimization (ERM) [20], it is paradoxical to depend on deep generative models that require large amounts of data to address data scarcity in subsequent tasks. This situation highlights the necessity to explore alternative approaches that can address the challenges of data-intensive models, utilizing a concept known as Vicinal Risk Minimization (VRM). VRM is a training principle that improves generalization in machine learning models by considering the vicinity of the training samples, hypothesizing that slight perturbations around real data points can yield more robust and better-performing models [21]. This concept seamlessly leads to the notion of data augmentation, a widely-adopted strategy in deep learning, especially in fields like computer vision where common methods of data augmentation include random cropping, rotation, and brightness adjustments [22]. With the emergence of one-shot generative models, a fresh motivation surfaces; harnessing their potential to generate realistic variations of training samples might offer a more sophisticated approach to VRM. Leveraging such generative capacities could potentially revolutionize the landscape of data augmentation and VRM practices. Recent one-shot generative models [23, 24, 25, 26], and their non-deep successors [27, 28, 29], generate patch-similar images and videos from a single target. Considering the patch-similarity as a vicinity criterion, one may therefore explore the ability of one-shot generative models for augmenting an entire dataset inspired by the VRM principle. In this study, our focus is on the application of one-shot generative models to sequences, which serve as a universal structure for ordered data, extending from images to time series. We aim to establish theoretical limits on their ability to augment datasets. Our work utilizes the Generating by Patch Distribution Matching (GPDM) algorithm, detailed in [28]. GPDM is designed to minimize the Wasserstein distance between patches of both real and generated images. Furthermore, this algorithm employs a multi-scale hierarchical strategy, facilitating an early understanding of the global image structure during the initial training stages. GPDM has emerged as the most efficient one-shot generative model capable of generating diverse samples, primarily due to its optimization based on Sliced Wasserstein distance (SWD). In contrast, other one-shot generative models either lack speed or do not employ metric distances, potentially compromising sample diversity [28]. For instance, Generative Patch Nearest Neighbors (GPNN)  [27] utilizes K-Nearest Neighbors (KNN), a similarity measure that is non-metric and results in less diverse samples compared to GPDM. Additionally, deep models like SinGAN [23] require extensive training times for each sample, which is inefficient compared to the model-free approach of GPDM.
@@ 3 Sequence and subsequence similarity
According to the definition, $\Omega_{S}$ is the set of all matrices created by choosing $d^{\prime}$ rows from the identity matrix $\mathbf{I}_{d}$. The probability distribution for singleton events in $\tilde{\Omega}_{S}\subseteq\Omega_{S}$ is uniform. Furthermore, the subset $\tilde{\Omega}_{S}$ must be sufficiently large to allow for the recovery of the sequence $\mathbf{X}$ from the projections $(\mathbf{s}^{T}\mathbf{s}\mathbf{X})$, as per (2). This suggests that $\bm{\lambda}\in\mathbb{N}^{d}$, since $\bm{\lambda}$ represents the frequency of occurrences of elements from $\mathbf{X}$ in these projections, as illustrated in Fig. 1. Based on this definition, we can proceed to define the behavior of a one-shot generative model. This model is designed to produce subsequence-similar sequences for every target sequence within the dataset. Before proceeding, we elucidate our understanding of similarity by referencing a conventional metric distance used to compare probability distributions. The Wasserstein distance is a powerful metric for quantifying the difference between two probability distributions. This metric, often referred to as the earth mover’s distance, provides an intuitive measure of the effort required to transform one distribution into another by considering the cost of moving distribution mass. Unlike other distance metrics that measure differences at fixed points, the Wasserstein distance considers the geometry of the space by accounting for the movement of distribution mass across it. To facilitate a deeper understanding, we will break down the mathematical formulation of the Wasserstein distance.
@@
$\begin{split}&\|\mathbf{X}^{\prime}-G(\mathbf{X},\mathbf{Z})\|\\ \leq&\textstyle\bm{\lambda}_{\min}^{-1}\sum_{\mathbf{s}\in\tilde{\Omega}_{S}}\norm{\mathbf{s}^{T}\mathbf{sX^{\prime}}-\mathbf{s}^{T}\mathbf{s}G(\mathbf{X},\mathbf{Z})}\\ =&\textstyle|\tilde{\Omega}_{S}|\bm{\lambda}_{\min}^{-1}\mathbb{E}_{S}\norm{\mathbf{S}^{T}\mathbf{SX^{\prime}}-\mathbf{S}^{T}\mathbf{S}G(\mathbf{X},\mathbf{Z})}\end{split}$ (14)
@@
Assume that $\forall\delta>0$, there exists a measurable function $G$ such that: $\mathbb{E}_{X}\mathbb{E}_{Z}\inf\mathbb{E}_{S,S^{\prime}}\norm{\mathbf{S}\mathbf{X}-\mathbf{S}^{\prime}G(\mathbf{X},\mathbf{Z})}\leq\delta$. Let $\mathbf{X}=\textup{vec}(\mathbf{Y})$ where $\mathbf{Y}$ belongs to $(\Omega_{Y}\subset\mathbb{R}^{n\times n},P_{Y})$. Then, if $\tilde{\Omega}_{S}$ denotes the set of all $n^{\prime}\times n^{\prime}$ substrings where $n^{\prime}\leq n$, we have following bounds:
@@
Remark 1.
In the preceding corollary, we presented results for square signal dimensions, i.e. $n\times n$, to maintain expositional clarity and enhance the elegance of the derivations. The bounds discussed can be readily extended to accommodate rectangular signal dimensions, such as $m\times n$ which involves analogous derivations.
4 Experimental setup
@@ 4.1 Dataset
RF sensing techniques for UAV detection and identification are practical for real-world use. These methods do not depend on UAVs’ wireless protocols like Bluetooth, 4G, or WiFi, and can manage physical differences across UAV models. This study utilizes a UAV RF dataset [7, 30] which encompasses RF signals obtained from three distinct UAV variants across a range of flight scenarios and is readily available for public access [^1] https://github.com/Al-Sad/DroneRF. Specifically, the dataset supports three primary tasks: the detection of a UAV’s presence (two classes), the detection of a UAV’s presence coupled with the identification of its type (four classes), and the detection of a UAV’s presence alongside the identification of its type and flight mode (ten classes). In this work, we skip the binary detection task as it may not be a challenge even for limited RF environments. For comprehensive details regarding the experimental setup—including the types of UAVs, the characteristics of the RF environments, and the data collection methodologies—readers are directed to the original publication. Addressing these aspects is beyond the scope of this work, which is focused on enhancing the performance of UAV identification using existing datasets. To facilitate the objectives of UAVs detection and identification, the authors employed Fourier transform to process the RF signals, thereby extracting their frequency components. RF signals are processed using Discrete Fourier Transform (DFT) for several key reasons. DFT reveals important frequency-domain characteristics often hidden in time-domain signals, reduces data dimensionality, and ensures uniform signal length, which is particularly beneficial when raw RF data may vary due to different UAV flight times. Furthermore, DFT facilitates the extraction of specific frequency features crucial for classification tasks and produces features that are generally more robust against common issues like Electromagnetic Interference (EMI) and Radio Frequency Interference (RFI) in UAV RF signals. Therefore, classifiers tend to perform better when trained on DFT-based features due to their inherent structure and discrimination power, leading to improved accuracy and generalization. The DFT-transformed RF signals in the referenced dataset consist of 2048 frequency features. For our application, we select the first 2025 features and organize them into a $45\times 45$ matrix structure to facilitate one-shot generation. This specific choice serves to validate the theoretical bounds presented in Corollary 2. As noted in Remark 1, adopting a more general approach for rectangular dimensions, such as reshaping the signal into a $32\times 64$ matrix, is also feasible. Additionally, omitting the last 23 frequency features from the total of 2048 has a negligible impact on the performance of the identification task. We also randomly selected 1,135 samples out of 22,700 (5%) from the dataset to simulate conditions of a low-data regime, ensuring data integrity and preserving the proportions of labels. Each sample in this dataset is associated with a label, which is a one-hot encoded vector sized either four or ten, depending on the classification task, with the first class designated for background noise in no-UAV scenarios. Initially, we apply a K-fold strategy (with $K=5$) to the original dataset, setting aside a consistent 20% random sample for testing and verification in each fold, with proportional representation of each label. This method ensures that these test samples remain unused in the training phase, preserving their utility for evaluating model performance against adversarial samples. The training phase then exclusively employs the remaining 80% of the dataset, totaling 908 samples in each fold which is called “complete” data throughout this work. For a comprehensive analysis, we further segment the training portion of each fold into reduced datasets, representing five scenarios with sizes between $10\sim 50\%$ of the training data. This segmentation allows us to demonstrate our model’s effectiveness across a range of reduced training dataset sizes. To ensure fair model performance comparisons, 5000 synthetic samples are created from subsampled sets for all generative models. This method facilitates meaningful evaluations of models on synthetic versus original data across varying scenarios.